The random variable Y = LogX has N(10,4) distribution. Find (a) The pdf of X (b) Mean and variance of X (c) P ( X ≤ 1000


1
Expert's answer
2021-11-15T15:48:05-0500

for distribution N(10,4):

"\mu=10, \sigma^2=4"


a)

for pdf of Y:


"f_Y(y)=\frac{1}{\sigma\sqrt{2\pi}}e^{-\frac{1}{2}(\frac{y-\mu}{\sigma})^2}=\frac{1}{2\sqrt{2\pi}}e^{-\frac{1}{2}(\frac{y-10}{2})^2}"


"\int^{\infin}_{-\infin}f_Y(y)dy=1"


"dy=e^{-y}dx"


then:


"f_Y(y)=e^yf_{X>0}(x)=2e^yf_{X}(x)"


"f_{X}(x)=f_Y(y)/(2e^y)=\frac{1}{4x\sqrt{2\pi}}e^{-\frac{1}{2}(\frac{logx-10}{2})^2}"


b)

mean of X:


"E(X)=\int^{\infin}_0xf_X(x)dx=\frac{1}{4\sqrt{2\pi}}\int^{\infin}_0e^{-\frac{1}{2}(\frac{logx-10}{2})^2}dx="


"=\frac{e^{12}}{4}erf(\frac{lnx-14}{2\sqrt2})|^{\infin}_0=e^{12}/4"


where erf is the error function.


variance of X:

"Var(X)=E(X^2)-(E(X))^2"


"E(X^2)=\int^{\infin}_0x^2f(x)dx=\frac{1}{4\sqrt{2\pi}}\int^{\infin}_0xe^{-\frac{1}{2}(\frac{logx-10}{2})^2}dx="


"=\frac{e^{28}}{4}erf(\frac{lnx-18}{2\sqrt2})|^{\infin}_0=e^{28}/4"


"Var(X)=e^{28}\/4-e^{24}\/4"


c)

"P ( X ≤ 1000)=\int^{1000}_0f(x)dx=\frac{1}{4\sqrt{2\pi}}\int^{1000}_0\frac{e^{-\frac{1}{2}(\frac{logx-10}{2})^2}}{x}dx="


"=\frac{1}{4}erf(\frac{lnx-10}{2\sqrt2})|^{1000}_0=\frac{1}{4}(erf(\frac{ln1000-10}{2\sqrt2})+1)=0.0765"